OpenStudy (anonymous):

PLEASE HELP ME!! Which statement best explains whether the table represents a linear or nonlinear function? Input (x) Output (y) −2 7 −5 4 −8 1 −11 −2 It is a linear function because there is a constant rate of change in both the input and output values. It is a nonlinear function because there is a constant rate of change in both the input and output values. It is a linear function because the input values are decreasing. It is a nonlinear function because the input values are increasing.

OpenStudy (anonymous):

@Frostbite

OpenStudy (anonymous):

ill help u

OpenStudy (anonymous):

ok thank you so much which one do you think it is

OpenStudy (anonymous):

what is this on a test or something

OpenStudy (anonymous):

???

OpenStudy (anonymous):

A linear equation is, basically, a line. An example is y=2x+3. They are commonly found in the form y=mx+b where m is the slope, and b is the y-intercept (where the line crosses the y-axis). A non-linear equation can be many things. Parabolas and exponential equations are not linear.

OpenStudy (anonymous):

sorry its this

OpenStudy (australopithecus):

You can either graph the points and if you get a straight line it is linear, if it deviates at all from a straight line it is non-linear

OpenStudy (australopithecus):

A linear function has a constant slope hence the equation y = mx + b, m and b are constants where m is the slope and b is the y-intercept. A non-linear function has a slope that changes (is not constant). The only condition for a function to be linear is that it has to have a constant slope it doesnt matter if the slope is negative or positive as long as it is constant (not changing). You can use the formula, \[slope\ = m\ =\frac{y_2 - y_1}{x_2 - x_1}\] you would have to systematically check in between each two point, if you get the same number out each time you have a linear equation

OpenStudy (australopithecus):

Another option is just to try to construct a linear equation out of two of the points by using the equation above to find the slope between two point. Then use slope in the formula y = mx + b sub one of the points in used to determine the slope and determine b. Then lastely plug in all x values given to see if you get the same y values out

OpenStudy (australopithecus):

If you had a table like this X | Y 1 | 2 3 | 4 5 | 3 a point is denoted as (x, y) use the slope formula with points (1,2) and (3,4) m = (4 - 2)/(3 - 1) = 2/2 = 1 thus we plug it into y=mx + b, we get y = 1*x + b now to find b we use (1,2) or we could use (3,4) 2 = 1*1 + b b = 2 - 1 b= 1 therefore the y intercept is 1 so we have the equation y = x + 1 now we test other points (5, 3), so we plug 5 into x to see if we get 3 out as a y value if we dont then the equation is non-linear y = 5 + 1 y = 6 thus the equation is not linear because 6 does not equal 3

OpenStudy (australopithecus):

If you get a different slope or the equation you set up fails to predict a y value in your table then it is non-linear.